Can someone tell me if I am going the right way?
Determine the correct ratio of y to x.
7/3y=5/4x=15/28
5/7y=10/21x=3/2
PLEASE TELL ME IF I DID THIS RIGHT OR DO I HAVE TO REVERSE THE RATIOS.....

Answer by Alan3354(21583)

Determine the correct ratio of y to x.
7/3y=5/4x=15/28
5/7y=10/21x=3/2
---------------
7y/3 = 5x/4
28y = 15x
y/x = 15/28
--------------
5y/7 = 10x/21
105y = 70x
y/x = 70/105
y/x = 2/3